sunisround:

Find k so that the quotient (x^4-8x^3+kx^2+15x-18)/x-6 has a zero remainder?

Vocaloid:

let x = 6 and find a value that makes the entire function 0

sunisround:

so by synthetic division?

Vocaloid:

just algebraic substitution

sunisround:

how do i do that?

Vocaloid:

replace "x" with 6 and set the function equal to 0 then find k

Vocaloid:

[just the numerator]

sunisround:

let me try it.

sunisround:

so k is equal to 10?

Vocaloid:

yup that's your answer

sunisround:

so that will give me a zero remainder?

Vocaloid:

yes

sunisround:

ok thank you again.

sillybilly123:

I dunno The problem is to find k so that: \( \dfrac{x^4-8x^3+kx^2+15x-18}{x-6} = \alpha , ~~~ \alpha \in \mathbb Z \) Or: \( (x^4-8x^3+kx^2+15x-18) = \alpha (x-6), ~~~ \alpha \in \mathbb Z \) Whichever way, solving by setting x = 6 and solving the numerator for 0 amounts to dividing 0 by 0 So specify that \(x \ne 6\) but find any other value of x that sets the numerator to zero works Yes ?!?!

Vocaloid:

I noticed that too, I will admit it's not the most mathematically sound method I could have used

sillybilly123:

😂